# GET-Zero: Graph Embodiment Transformer for Zero-shot Embodiment Generalization

### v1

@@ Abstract
Keywords: Cross-embodiment learning, Dexterous manipulation
1 Introduction
@@
Motivated by this limitation, the goal of this work is to introduce an embodiment representation that explicitly leverages the graph connectivity of the robot to improve zero-shot control of robots with variations in graph structure. We propose GET-Zero, which introduces the Graph Embodiment Transformer (GET) model architecture that enables zero-shot adaptation to new embodiments (Fig. 1). The key idea of GET is to leverage the robot graph connectivity as a structural bias in the transformer attention mechanism. By training this model using behavior data from a variety of robot embodiments, we can then zero-shot control previously unseen embodiments with varied geometry, number of joints, and graph structure. Specifically, GET-Zero consist of three novel components:
@@
    Embodiment-Aware Distillation. To train GET-Zero we first collect demonstration data from embodiment-specific experts. By conditioning on embodiment information, we then distill knowledge from embodiment-specific experts into an embodiment-aware GET model using behavior cloning (BC) that controls both training and unseen embodiments with a single set of network weights. Our approach contrasts with prior methods [2, 3, 4] that use RL to simultaneously learn an embodiment-aware model while learning to complete the task across all embodiments, which is significantly more difficult and prevents the use of expert demonstration data.
@@
    Self Modeling Loss. During the distillation BC phase, we introduce an additional self modeling loss alongside the standard policy action loss. The self-modeling output predicts the position of each joint in 3D space (i.e., forward kinematics), which encourages the network to understanding the embodiment graph and infer actions accordingly. We found that this simple and general supervision improves generalization performance.
@@ 2 Related Work
Cross-Embodiment pretraining with embodiment-specific finetuning. Recent approaches leverage large-scale pretraining, either on visual data, robot trajectories, language, or tasks as an initialization for policy learning or generalist agents [9, 10, 11, 12, 13, 14, 15]. Other approaches leverage robot data across embodiments to initialize a generalist agent, and then fine-tune the base model or an action head to adapt to a new robot embodiments [16, 17]. These method often assume a unified action space, which is reasonable for navigation [18] or for robot arm policies where the end-effector pose is a shared representation across arms [17, 19, 16, 20]. However, unified actions spaces are often not sufficient for tasks where precise, low-level joint actions matter, such as dexterous manipulation [21, 22, 23, 24]. Moreover, many approaches still require non-trivial amounts of additional training data specific to the target hardware and environment. Instead, our method uses the graph structure of the robot as a general embodiment representation to improve zero-shot transfer for tasks where a unified action space is not possible.
@@
Embodiment Representation. Prior works in multi-embodiment policies demonstrate using a single policy to control different robot designs [1, 25, 2, 3, 4, 6, 5]. However, most methods show limited generalization ability to new embodiment designs beyond the training set due to incomplete embodiment representations. For example, Octo [1] has 3rd-person images of the robot arm which doesn’t generalize across robot appearance changes, though this could be addressed with robot inpainting [26]. Other methods learn cross embodiments representations that align actions across different embodiments [27, 28, 29]. In contrast, GET-Zero uses the graph embodiment structure connecting local sensory information as a flexible and general representation across robots.
@@
Embodiment-aware policy architectures. Most related to our approach is the line of work developing embodiment-aware policies. Prior methods [30, 31, 2] structure a graph neural network (GNN) to match the embodiment graph with joints as nodes and links as edges to control stick-like walking characters in simulation. Kurin et al. [3] shows that a GNN matching the embodiment graph is outperformed with the fully connected attention mechanism in transformers [7] despite having no graph encoding. Gupta et al. [4] extends this model with dynamics and kinematics encodings, demonstrating zero-shot generalization to these properties. However, these methods do not generalize well to unseen embodiment graphs due to no [3] or limited [4] graph representations. Our method GET-Zero extends the transformer architecture in [3, 4] to include an explicit graph representation that improves zero-shot generalization to unseen embodiment graphs. Furthermore, unlike prior embodiment-aware RL methods that do locomotion only in simulation [30, 31, 2, 3, 4], GET-Zero uses behavior cloning from experts for an in-hand rotation task transferable to a real robot.
3 Method: Graph Embodiment Transformer
@@
We present Graph Embodiment Transformer (GET), an embodiment-aware transformer encoder [7] that uses the embodiment graph as a structural bias in network architecture (Fig. 2). An embodiment graph consists of nodes representing local joint information and directed edges representing links connecting parent and child joints. GET represents each joint as separate transformer tokens containing both local sensory observations and local hardware properties (§3.1). The graph edges (i.e., links) are encoded through a learned attention bias in the self-attention layers (§3.2).
3.1 Embodiment Tokenization
@@
Observations can either be local $l$ to specific joints or global $g$ across the entire embodiment. Additionally, observations are either fixed $f$ or variable $v$ if they change during execution. This yields four observation types: variable local $o_{vl}$ (e.g., joint angle/velocity), variable global $o_{vg}$ (e.g., a global time encoding, or environment state), fixed local $o_{fl}$ (e.g., joint position in rest pose, or joint limit ranges) and fixed global $o_{fg}$ (e.g., a task identifier). $H$ past history steps are included only for variable observations ($o_{vl}$ and $o_{vg}$). The local observations for joint $j$ are defined as $o_{*l,j}$. The transformer token $T_{j}$ for joint $j$ at timestep $t$ is constructed as $T_{j}^{t}=[o_{vg}^{t\rightarrow t-H},o_{fg},o_{vl,j}^{t\rightarrow t-H},o_{fl,j}]$. The tokens pass through a learned linear embedding before entering the transformer encoder.
@@
3.2 Graph Encoding
One challenge is that the transformer encoder has no direct graph encoding mechanism in the original implementation [7]. Input tokens are permutation invariant without a positional encoding mechanism, but are often linearly encoded with sinusoidal or learned positional encodings. Prior embodiment-aware methods address this in multiple ways. Kurin et al. [3] use no positional encoding meaning no graph representation is present. Gupta et al. [4] linearize the graph using a depth-first search ordering then apply a learned linear positional encoding. However, this approach is sensitive to graph variations as the DFS ordering is not unique, which empirically caused a $\sim 75\%$ drop in performance with the opposite node order in Gupta et al. [4]. Sferrazza et al. [32] use the adjacency matrix as a binary attention mask to limit communication early in the transformer to adjacent nodes.
Our GET model uses a learned attention bias in the encoder self-attention mechanism to explicitly encode the graph similar to the Graphormer work by Ying et al. [33]. Using the notation from Ying et al. [33], the inputs to the self-attention layer given $J$ joints are the hidden states $H=[h_{1}^{\top},...,h_{J}^{\top}]\in\mathbb{R}^{J\times d}$. $H$ is projected by three matrices $W_{Q}\in\mathbb{R}^{d\times d_{K}}$,$W_{K}\in\mathbb{R}^{d\times d_{K}}$, $W_{V}\in\mathbb{R}^{d\times d_{V}}$ to produce $Q=HW_{Q}$, $K=HW_{K}$, $V=HW_{V}$. Then the attention layers will compute an attention matrix $A\in\mathbb{R}^{J\times J}$ where $A_{ij}$ represents the attention score computed between joints $i$ and $j$ as follows: $A=QK^{\top}/\sqrt{d},\text{Attn}(H)=\text{softmax}(A)V$.
To encode the embodiment graph, we learn two separate biases to the attention scores ($A$): a Spatial Encoding [33], and a novel Parent-Child Encoding, detailed below.
Spatial Encoding. The spatial encoding computes the shortest path distance (SPD) $\phi^{\text{SPD}}$ between node $i$ and node $j$ in the graph as $\phi^{\text{SPD}}(i,j)$. An embedding table $s$ is indexed with the SPD distance as $s_{\phi^{\text{SPD}}(i,j)}$ to produce a learned scalar that is added on top of the attention score $A_{ij}$.
@@ 3.2 Graph Encoding
Parent-Child Encoding. Unlike Graphormer [33] that only encodes undirected graph in the attention mechanism, we introduce a new parent-child attention bias to encode directed features. For joints $i$ and $j$, we compute the parent distance $\phi^{P}(i,j)=\phi^{\text{SPD}}(i,j)\mathbbm{1}\{i\text{ is parent of }j\}$ which is the distance between $i$ and $j$ if $j$ is the child of $i$ at some distance along the forward kinematic chain, otherwise 00. We compute an analogous formula for the child distance as $\phi^{C}(i,j)=\phi^{\text{SPD}}(i,j)\mathbbm{1}\{i\text{ is child of }j\}$. There are associated scalar embedding tables $p$ and $c$.
The intuition is that the model may attempt to compute forward kinematics through the embodiment graph and the parent-child encoding enables the model to modulate attention based on directed nature of a forward chain (child joint angles do not impact the parent joint’s 3D pose).
Attention Bias Computation. For each attention head and encoder layer, we have separate embeddings $s$, $p$, and $c$ and compute attention as: $A_{ij}=\frac{(h_{i}W_{Q})(h_{j}W_{K})^{T}}{\sqrt{d}}+s_{\phi^{\text{SPD}}(i,j)}+p_{\phi^{\text{P}}(i,j)}+c_{\phi^{\text{C}}(i,j)}$.
@@
Output Heads. After passing $J$ tokens through each transformer encoder, GET produces $J$ corresponding latent features. We include a policy head to predict actions and a self-modeling head used to predict meta-features about the embodiment. Further details on these heads are discussed in §4.3.
4 Case Study: In-hand Object Rotation
We conduct a case study on applying GET-Zero to an in-hand object rotation task with different variations of LEAP Hand [8]. Unlike pick-and-place tasks where a unified action space across different robot arms is easy to define (e.g., end-effector space), the task of in-hand object rotation requires the policy to differentiate low-level joint space actions that are aware of the embodiment structure. Applying GET-Zero to the in-hand rotation task requires three stages (Fig. 3): 1) generate embodiment variations (§4.1), 2) train embodiment-specific experts using RL (§4.2), and 3) distill knowledge from experts into an embodiment-aware GET architecture (§4.3).
4.1 Procedural Embodiment Generation
@@
4.2 Train Embodiment-Specific Experts
We train embodiment-specific RL experts using PPO [34] in Isaac Gym [35] to complete the in-hand rotation task, adapted from the setup by Shaw et al. [8]. This setup consists of rewards for rotation along the yaw axis, penalties for high motor torques and dropping the cube, and domain randomization across cube sizes. These learned policies operate only on proprioceptive joint states and predict delta joint targets that guide a PD controller. We pick the best of five training seeds for each embodiment and filter out embodiments that are not able to complete a full $2\pi$ rotation within $30$s. Many hand designs are unable to complete the rotation task under this threshold due to many missing joints, resulting in 44 training embodiments with associated expert policies.
4.3 Distill Experts into Embodiment-Aware Transformer
@@
Separating the collection of demonstration data from the training of embodiment-aware policies gives more freedom to use any existing policy architecture or training setup to generate expert demonstrations (e.g. teleoperation, RL, optimization based methods, privilaged information, etc.). Prior embodiment-aware methods [3, 4] do not separate these stages and instead jointly learn the embodiment-aware policy along with learning how to control training embodiments.
GET Training Setup. We train the GET architecture (§3) using BC. Specifically, our token input consists normalized joint angles and PD target joint states ($o_{vl}$), 3D joint position and joint rotation with respect to the parent joint from the URDF file ($o_{fl}$), and a sinusoidal phase encoding with a period of two seconds to encode cyclic progression through the rotation cycle ($o_{vg}$). The directed embodiment graph is an additional input used in the graph attention encoding (§3.2).
Policy Head. The network contains a policy head for predicting single-step action. Each joint has a PD target state that is initialized to the starting joint state. The action head predicts delta actions that are added to this target state, which sets the target state for the low-level PD controller. We supervise these actions using $L_{2}$ loss with the demonstration data.
Self-Model Head. The network also has a self-modeling head to predict the current 3D pose of each joint in the robot local frame. This forward kinematics (FK) task is simple and generally applicable across embodiments, yet requires the network to use the spatial relationships from the graph and current joint angles along the FK chain. We supervise with the ground truth FK pose with $L_{2}$ loss. As the policy and self-modeling heads operate on the same latent representations output by GET, self-modeling encourages the actions to be predictable from an embodiment-rich representation.
5 Experiments
We procedurally generate 236 graph variations of the LEAP hand [8] by removing various combinations of joints and associated links from the embodiment (§4.1). From these, we obtain 44 embodiments with graph variations with associated expert RL policies that achieve a baseline level of performance (§4.2). These 44 embodiments serve as training embodiments and only have graph variations (link length extensions are not seen during training). Our experiments consist of using demonstration data from these embodiments to train various baselines and ablations of the GET architecture to evaluate zero-shot performance to unseen embodiment graphs and link extensions.

### v2

@@ Abstract
I Introduction
@@
Motivated by this limitation, the goal of this work is to introduce an embodiment representation that explicitly leverages the graph connectivity of the robot to improve zero-shot control of robots with variations in graph structure. We propose GET-Zero, which introduces the Graph Embodiment Transformer (GET) model architecture that enables zero-shot adaptation to new embodiments (Fig. 1). The key idea of GET is to leverage the robot graph connectivity as a structural bias in the transformer attention mechanism. By training this model using behavior data from a variety of robot embodiments, we can then zero-shot control previously unseen embodiments with varied geometry, number of joints, and graph structure. Specifically, GET-Zero consist of three components:
@@
    Embodiment-aware Distillation. To train GET-Zero we first collect demonstration data from embodiment-specific experts. By conditioning on embodiment information, we then distill knowledge from embodiment-specific experts into an embodiment-aware GET model using behavior cloning (BC) that controls both training and unseen embodiments with a single set of network weights. Prior methods [2, 3, 4] use RL to simultaneously learn an embodiment-aware model while learning to complete the task across all embodiments. In contrast, our method simplifies data generation by independently learning experts, then jointly distilling expert behavior.
@@
    Self-modeling Loss. During the distillation BC phase, we introduce a self-modeling loss to predict the position of each joint in 3D space (i.e., forward kinematics). We found that this simple and general supervision improves zero-shot performance. Self-supervision is common in NLP [8] and vision [9] domains, but is less explored in cross-embodiment learning [10, 11, 12].
@@ II Related Work
Cross-Embodiment pretraining with embodiment-specific finetuning. Recent approaches leverage large-scale pretraining, either on visual data, robot trajectories, language, or tasks as an initialization for policy learning or generalist agents [14, 15, 16, 17, 18, 19, 20]. Other approaches leverage cross-embodiment robot data to initialize a generalist agent, and then fine-tune the base model or an action head to adapt to new robot embodiments [21, 22]. These method often assume a unified action space, which is reasonable for navigation [23] or for robot arm policies where the end-effector pose is a shared representation across arms [22, 24, 21, 25]. However, unified actions spaces are often not sufficient for tasks where precise, low-level joint actions matter, such as dexterous manipulation [26, 27, 28, 29]. It is challenging to formulate a unified action space in the case when hands have varying number of fingers and kinematic feasibility and when contact may occur anywhere along the fingers. Instead, our method uses the robot graph structure as an embodiment representation to perform joint-level control for tasks where a unified action space is not possible. GET-Zeroalso does not assume demonstration data on the target embodiment is available nor require finetuning.
@@
Embodiment Representation. Prior multi-embodiment policies demonstrate using a single policy to control different robot designs [1, 30, 2, 3, 4, 6, 5]. However, most methods show limited generalization ability to new embodiment designs beyond the training set due to incomplete embodiment representations. For example, Octo [1] has 3rd-person images of the robot arm which doesn’t generalize across robot appearance changes, though this could be addressed with robot inpainting [31]. Other methods learn cross-embodiment representations that align actions across different embodiments [12, 32, 11]. In contrast, GET-Zerouses an embodiment graph composed of joints which is readily accessible from URDF files as a practical representation when joint-level actions matter.
@@
Embodiment-aware policy architectures. Most related to our approach is the line of work developing embodiment-aware policies. Prior methods [33, 34, 2] structure a graph neural network (GNN) to match the embodiment graph with joints as nodes and links as edges to control stick-like walking characters in simulation. [3] shows that a GNN matching the embodiment graph is outperformed with the fully connected attention mechanism in transformers [7] despite having no graph encoding. [4] extends this model with dynamics and kinematics encodings, demonstrating zero-shot generalization to these properties. However, these methods do not generalize well to unseen embodiment graphs due to no [3] or limited [4] graph representations. Our method GET-Zero extends the transformer architecture in [3, 4] to include an explicit graph representation that improves zero-shot generalization to unseen embodiment graphs. Furthermore, unlike prior embodiment-aware RL methods that do locomotion only in simulation [33, 34, 2, 3, 4], GET-Zero uses behavior cloning from experts for an in-hand rotation task transferable to a real robot.
III Method: Graph Embodiment Transformer
@@
We present Graph Embodiment Transformer (GET), an embodiment-aware transformer encoder [7] that uses the embodiment graph as a structural bias in network architecture (Fig. 2). An embodiment graph consists of nodes representing local joint information, directed edges representing links connecting parent and child joints, and undirected edges between joints at the start of independent serial chains. GET represents each joint as separate transformer tokens containing both local sensory observations and local hardware properties (§III-A). The graph edges (i.e., links) are encoded through a learned attention bias in the self-attention layers (§III-B).
III-A Embodiment Tokenization
@@
Observations can either be local $l$ to specific joints or global $g$ across the entire embodiment. Additionally, observations are either fixed $f$ or variable $v$ if they change during execution. This yields four observation types: variable local $o_{vl}$ (e.g., joint angle/velocity), variable global $o_{vg}$ (e.g., a global time encoding, or environment state), fixed local $o_{fl}$ (e.g., joint position in rest pose, or joint limit ranges) and fixed global $o_{fg}$ (e.g., a task identifier). $H$ past history steps, indicated $t\rightarrow t-H$, are included for variable observations. The local observations for joint $j$ are defined as $o_{*l,j}$. The transformer token $T_{j}$ for joint $j$ at timestep $t$ is constructed as $T_{j}^{t}=[o_{vg}^{t\rightarrow t-H},o_{fg},o_{vl,j}^{t\rightarrow t-H},o_{fl,j}]$. The tokens pass through a learned linear embedding before entering the transformer encoder.
@@
III-B Graph Encoding
One challenge is that the transformer encoder has no direct graph encoding mechanism in the original implementation [7]. Input tokens are permutation invariant without a positional encoding mechanism, but are often linearly encoded with sinusoidal or learned positional encodings. Prior embodiment-aware methods address this in multiple ways. [3] use no positional encoding meaning no graph representation is present. [4] linearizes the graph using a depth-first search ordering then apply a learned linear positional encoding. However, this approach is sensitive to graph variations as the DFS ordering is not unique, which empirically caused a $\sim 75\%$ drop in performance with the opposite node order in [4]. [35] use the adjacency matrix as a binary attention mask to limit communication early in the transformer to adjacent nodes.
Our GET model uses a learned attention bias in the encoder self-attention mechanism to explicitly encode the graph similar to the Graphormer work by [36]. We opt for Graphormer for its simple implementation, but note that our approach would benefit as graph transformers improve [37, 38]. Using the notation from [36], the inputs to the self-attention layer given $J$ joints are the hidden states $H=[h_{1}^{\top},...,h_{J}^{\top}]\in\mathbb{R}^{J\times d}$. $H$ is projected by three matrices $W_{Q}\in\mathbb{R}^{d\times d_{K}}$,$W_{K}\in\mathbb{R}^{d\times d_{K}}$, $W_{V}\in\mathbb{R}^{d\times d_{V}}$ to produce $Q=HW_{Q}$, $K=HW_{K}$, $V=HW_{V}$. Then the attention layers will compute an attention matrix $A\in\mathbb{R}^{J\times J}$ where $A_{ij}$ represents the attention score computed between joints $i$ and $j$ as follows: $A=QK^{\top}/\sqrt{d},\text{Attn}(H)=\text{softmax}(A)V$.
To encode the embodiment graph, we learn two separate biases to the attention scores ($A$):
Spatial Encoding. The spatial encoding computes the shortest path distance (SPD) $\phi^{\text{SPD}}$ between node $i$ and node $j$ as $\phi^{\text{SPD}}(i,j)$, treating all edges as undirected. An embedding table $s$ is indexed with the SPD distance as $s_{\phi^{\text{SPD}}(i,j)}$ to get a learned scalar that is added to the attention score $A_{ij}$.
@@ III-B Graph Encoding
Parent-Child Encoding. Unlike Graphormer [36] that only encodes undirected graph in the attention mechanism, we introduce a new parent-child attention bias to encode directed features. For joints $i$ and $j$, we compute the parent distance $\phi^{P}(i,j)=\phi^{\text{SPD}}(i,j)\mathds{1}\{i\text{ is parent of }j\}$ which is the distance between $i$ and $j$ if $j$ is the child of $i$ at some distance along the forward kinematic chain, otherwise 00. We compute an analogous formula for the child distance as $\phi^{C}(i,j)=\phi^{\text{SPD}}(i,j)\mathds{1}\{i\text{ is child of }j\}$. There are associated scalar embedding tables $p$ and $c$. This means that the directed edges have distinct attention biases if we go from parent-to-child or child-to-parent.
Our motivation is that the parent-child encoding lets the model modulate attention based on directed connectivity of the forward joint chain, which could be helpful to compute forward kinematics.
Attention Bias Computation. For each attention head and encoder layer, we add attention bias embeddings $s$, $p$, and $c$ to compute attention as: $A_{ij}=\frac{(h_{i}W_{Q})(h_{j}W_{K})^{T}}{\sqrt{d}}+s_{\phi^{\text{SPD}}(i,j)}+p_{\phi^{\text{P}}(i,j)}+c_{\phi^{\text{C}}(i,j)}$.
@@
Output Heads. After passing $J$ tokens through the transformer encoder, GET produces $J$ corresponding features. We include a policy head to predict actions and a self-modeling head used to predict meta-features about the embodiment. More details on the heads are discussed in §IV-C.
IV Case Study: In-hand Object Rotation
We conduct a case study on applying GET-Zero to in-hand object rotation with different variations of LEAP Hand [13]. Unlike pick-and-place tasks where a unified action space across different robot arms is easily defined (e.g., end-effector space), in-hand object rotation requires the policy to differentiate low-level joint space actions that are aware of the embodiment structure. Applying GET-Zero to in-hand rotation requires three stages (Fig. 3): 1) generate embodiment variations (§IV-A), 2) train RL embodiment-specific experts (§IV-B), and 3) distill knowledge from experts into an embodiment-aware GET architecture (§IV-C).
IV-A Procedural Embodiment Generation
@@
IV-B Train Embodiment-Specific Experts
We train embodiment-specific RL experts using PPO [39] in Isaac Gym [40] to complete the in-hand rotation task, adapted from the setup in [13]. This setup has rewards for rotation along the yaw axis, penalties for high motor torques and dropping the cube, and domain randomization across cube sizes. These learned policies operate only on proprioceptive joint states and predict delta joint targets that guide a PD controller. We pick the best of five RL seeds per embodiment and filter embodiments that do not complete a full $2\pi$ rotation within $30$s. Many hand designs can not complete the task under this threshold due to many missing joints, resulting in 44 training embodiments with associated expert policies. In general, it’s not necessary to train experts for the entire design space, which might be large, but instead on a subset of representative embodiments (see Fig. 5).
IV-C Distill Experts into Embodiment-Aware Transformer
@@
Separating collection of demonstration data from training embodiment-aware policies gives freedom to use existing policy architecture or training setups to generate expert demonstrations (e.g., teleoperation, RL, optimization methods, privileged information, etc.). Prior embodiment-aware methods [3, 4] instead jointly learn an embodiment-aware policy along with learning how to control training embodiments. This means the baselines do not support reusing existing demonstration data and require advanced replay buffer balancing as embodiments learn at different rates [4].
GET Training Setup. We train the GET architecture (§III) using BC. Specifically, our token input has normalized joint angles and PD target joint states ($o_{vl}$), 3D joint position and joint rotation with respect to the parent joint from the URDF file ($o_{fl}$), and a sinusoidal phase encoding with a period of two seconds to encode cyclic progression of the rotation cycle ($o_{vg}$). The directed embodiment graph is an additional input used in the graph attention encoding (§III-B).
Policy Head. A policy head predicts per-joint single-step action. Each joint has a PD target state that is initialized to the starting joint state. The action head predicts delta actions that are added to this target state, which sets the target state for the low-level PD controller. We supervise these actions using $L_{2}$ loss with the demonstration data.
Self-Model Head. A self-modeling head predicts the current 3D pose of each joint in the robot local frame. This forward kinematics (FK) task is simple and generally applicable across embodiments, yet requires the network to use the spatial relationships from the graph and current joint angles along the FK chain. We supervise with the ground truth FK pose with $L_{2}$ loss. Both the policy and self-modeling heads operate on the same output latent representations to encourage actions to be predictable from an embodiment-rich representation. We opt for per-joint FK for its simplicity, but note that more complex self-modeling is possible [41].
V Experiments
We procedurally generate 236 graph variations of the LEAP hand [13] by removing various combinations of joints and associated links from the embodiment (§IV-A). From these, we obtain 44 embodiments with graph variations with associated expert RL policies that achieve a baseline level of performance (§IV-B). These 44 embodiments serve as training embodiments and only have graph variations (link length extensions are not seen during training). Our experiments consist of using demonstration data from these embodiments to train various baselines and ablations of the GET architecture to evaluate zero-shot performance to unseen embodiment graphs and link extensions.